OpenStudy (anonymous):

Use a transformation to solve the equation. A. p = –36 B. p = –4 C. p = 4 D. p = 36

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{P}{-12}=-\frac{1}{3}\iff P=(-12)\times \left(-\frac{1}{3}\right)\]